# Rishanth Rajendhran1  Amir Zadeh2  Matthew Sarte2  Chuan Li2  Mohit Iyyer1

### v2

@@ Abstract
Metrics like FactScore and VeriScore that evaluate long-form factuality operate by decomposing an input response into atomic claims and then individually verifying each claim. While effective and interpretable, these methods incur numerous LLM calls and can take upwards of 100 seconds to evaluate a single response, limiting their practicality in large-scale evaluation and training scenarios. To address this, we propose VeriFastScore, which leverages synthetic data to fine-tune Llama3.1 8B for simultaneously extracting and verifying all verifiable claims within a given text based on evidence from Google Search. We show that this task cannot be solved via few-shot prompting with closed LLMs due to its complexity: the model receives $\sim$4K tokens of evidence on average and needs to concurrently decompose claims, judge their verifiability, and verify them against noisy evidence. However, our fine-tuned VeriFastScore model demonstrates strong correlation with the original VeriScore pipeline at both the example level ($r=\mathbf{0.80}$) and system level ($r=\mathbf{0.94}$) while achieving an overall speedup of 6.6$\times$ (9.9$\times$ excluding evidence retrieval) over VeriScore. To facilitate future factuality research, we publicly release our VeriFastScore model and synthetic datasets. [^1] Code and data available at https://github.com/RishanthRajendhran/VeriFastScore.
@@
0.18ex
[…]
Rishanth Rajendhran1  Amir Zadeh2  Matthew Sarte2  Chuan Li2  Mohit Iyyer1 1University of Maryland 2Lambda Labs {rishanth, miyyer}@umd.edu
@@ 1 Introduction
Despite these challenges, VeriFastScore matches the quality of VeriScore closely while offering substantial improvements in both cost and runtime efficiency: it achieves a Pearson correlation of $\mathbf{0.80}$ (Pearson r, p<0.001) with VeriScore’s scores, a 9.9$\times$ modeling speedup and a 6.64$\times$ overall speedup in wall-clock time, reducing both modeling and retrieval latency while maintaining the metric’s interpretability. Human evaluation corroborates the high quality of outputs from VeriFastScore. We release VeriFastScore and our synthetic training dataset to support efficient, interpretable factuality evaluation and facilitate its use in alignment settings.
@@ 2.3 VeriFastScore’s Integrated Approach
Despite the lack of decomposition at the retrieval stage, we find that VeriFastScore maintains strong performance. On average, each response contains $\sim$23 verifiable claims and is associated with $\sim$186 evidence snippets, totaling over 4k tokens. These are processed together, with VeriFastScore identifying and verifying each claim directly against the evidence context. As shown in Table 9 and Table 5, VeriFastScore achieves high factual precision and recall with respect to VeriScore, while reducing inference time per response by over 6$\times$.
@@ 2.4 Generating Synthetic Training Data for VeriFastScore
We selected prompts from the Tulu3 Personas dataset Lambert et al. (2025), which contains diverse and high-quality instructions across domains intended to test model’s instruction-following ability and adaptation to user intent. This dataset was chosen because it aligns well with our goal of generating model responses that contain factual, verifiable content. We used GPT-4o-mini to identify a subset of prompts likely to elicit factual claims [^6] More details in A.4. A manual inspection of 200 filtered prompts showed that over 90% did indeed lead to factually dense responses.
@@
claim-level GPT-4o few-shot ICL 0.30 0.31 18.1 7.2 73.2 0.28
@@
Table 3: Performance metrics across test splits: VeriFastScore produces claims and verification labels that are largely consistent with VeriScore, achieving a strong corelation of 0.80 even with sentence-level evidence, while significantly outperforming the GPT-4o few-shot prompting baseline. Evidence Granularity denotes whether evidence was retrieved using full sentences or claims: claim-level includes VeriScore data and Tulu3 Personas data while sentence-level includes Tulu3 Personas data only. Precision and Recall measure overlap between claims identified by VeriFastScore and VeriScore. Under Claim Accuracy, Correct, Incorrect (Label), and Missing (Claim) capture agreement, label mismatches, and omissions relative to VeriScore. A small fraction of instances were marked Erroneous (not shown) due to ambiguities in automatic evaluation (see A.5). Correlation reports Pearson $r$ between factuality scores (see Eq. 1) assigned by VeriFastScore and VeriScore. For all Pearson’s correlation scores reported in this paper, p<0.001, unless stated otherwise.
@@ 2.5 Evaluation Metrics
During early experiments, we found that exact match metrics—comparing model-extracted claims and verification labels directly to reference outputs—were often too brittle. Minor surface-level differences (e.g., paraphrasing, slight reordering, or omission of non-essential modifiers) led to false negatives, thereby underestimating the model’s actual performance. For example, VeriFastScore achieved an exact match claim accuracy of 23.7% on the test set, versus 71.6% with paraphrase-aware evaluation. [^8] We re-report Table 9 with exact match metrics in Appendix
@@ 3 Results
Response … Navigating a rainforest can pose several challenges … unpredictable weather which can lead to sudden rainstorms or flooding … VeriFastScore: (Supported) Unpredictable weather in a rainforest can lead to sudden rainstorms. VeriScore: (Unsupported) Navigating a rainforest can lead to sudden rainstorms.   In this example, VeriFastScore exhibits precise semantic alignment between extracted claim and model response while VeriScore doesn’t.
@@ 3.2 Speedup
We measured the execution time for both the VeriFastScore and VeriScore pipelines. By retrieving evidence at a coarser granularity, VeriFastScore achieves a 39% reduction in API costs and cuts retrieval time by almost half. Additionally, VeriScore takes significantly longer due to its sequential claim extraction and verification processes. VeriFastScore is approximately 10 times faster in extracting and verifying claims without sacrificing performance (see Table 9). Overall, VeriFastScore achieves a 6.64x speedup compared to VeriScore. Longer model responses further increase speedup, as VeriScore’s sentence-by-sentence processing is particularly slow.
@@ 3.3 Ranking factuality of models with VeriFastScore
As shown in Figure 2, factuality scores produced by VeriFastScore correlate strongly with those from VeriScore (Pearson $r=0.94$, p=$1.1e^{-5}$), indicating alignment despite differing approaches. Given that VeriFastScore operates over a single model pass and avoids per-claim retrieval, it delivers these rankings with significantly reduced latency and cost (see Table 5, 6).
@@ 4 Qualitative analysis
Compararive inspection of claims extracted by VeriFastScore and VeriScore shows four recurring missteps. First, the model sometimes stitches together unrelated clauses into unwarranted cause–effect claims. Second, it extracts statements using vague or anaphoric phrases that lack the context needed for standalone verification. Third, evidence retrieval using Google search can surface both supporting and contradicting passages for the same claim, causing label flips. And fourth, both approaches occasionally extract unverifiable claims from model response.
Human evaluation:
A manual review of 401 claims across 21 test instances found that VeriFastScore reliably grounded its outputs in the response text, with no hallucinated claims. Verification accuracy was high, though some false positives arose from noisy evidence. Common failure modes included missed claims (in 29% of examples) and unverifiable outputs due to missing context. Full details are in Appendix A.3.
@@ 6 Conclusion
Despite the complexity of the task, VeriFastScore closely matches the accuracy of VeriScore (Pearson $r=0.80$), while achieving a 6.6$\times$ speedup in evaluation time. It also outperforms strong few-shot prompting baselines such as GPT-4o, and provides interpretable, claim-level outputs suitable for evaluation and alignment.
@@ 7 Limitations
Our retrieval strategy issues Google search queries for each sentence in the model response, aggregating the results into a shared evidence context. While this aligns with the single-pass structure of VeriFastScore, it can introduce irrelevant or insufficient evidence, especially for vague or underspecified sentences. More adaptive or agentic retrieval—allowing the model to detect failures and issue improved follow-up queries—may yield higher evidence quality.
@@
VeriFastScore produces fine-grained factuality labels for claims but does not provide natural language justifications. Although rationale generation adds computational cost, it could improve interpretability and trustworthiness, especially in sensitive domains. Training with rationales, even if only used during finetuning (e.g., as in Qwen3’s “reasoning mode” Yang et al. (2025)), may offer benefits without incurring runtime overhead at inference time.
@@
We use GPT-4o-mini to assess label correctness and semantic equivalence, which enables scalable evaluation but introduces potential noise—particularly for nuanced paraphrase detection. This may affect metric estimates such as precision and recall, and obscure more subtle performance differences across models.
@@ A.1 Training VeriFastScore
As seen in Table 9, there was a $\sim$20drop in claim accuracy when using sentence-level evidence, which is our true test setting, instead of claim-level evidence. This is possibly due to differences in quality of retrieved search results and length of consolidated evidence depending on the granularity at which evidence was collected. To mitigate this, the model was further trained on a mixture of claim- and sentence-level evidence in the second stage of training.
@@
A.2 VeriFastScorer model
@@
VeriScore 20.5 23 \$0.01725
VeriFastScore 58.8 14 \$0.0105
A.3 Human Evaluation
Claim grounding and hallucination:
All extracted claims were grounded in the original model response, with no instances of hallucinated or fabricated claims. This is a crucial property for a factuality metric, particularly in downstream use cases like hallucination detection and RLHF training where spurious claim hallucination could mislead supervision.
Extraction coverage and omissions:
VeriFastScore was generally effective at identifying claims, especially in responses rich with factual content. However, it showed reduced reliability in responses that interweave factual and non-factual content, or contain few factual claims overall. Claim omission was observed in approximately 29% of evaluated instances, though two-thirds of these cases involved fewer than four missing claims. Notably, in 50% of omission cases, VeriFastScore extracted claims only from the final few sentences of the response, missing earlier factual content—suggesting a positional bias or early truncation.
Unverifiable and ill-formed claims:
Roughly 10% of the extracted claims were judged to be unverifiable due to missing key details (e.g., named entities, dates) in the original model response. In rare cases, VeriFastScore also extracted claims from inherently unverifiable sentences, such as advice or vague generalizations. Ill-formed claims—those missing critical contextual elements—were rare (around 1%), but they did alter the meaning of the original sentence in notable ways.
Verification accuracy:
VeriFastScore was largely accurate in labeling claims against the provided evidence, with an estimated verification error rate of 8.5%. Among these errors, 76% were false positives—instances where VeriFastScore incorrectly marked a claim as supported. These errors often arose when partial matches or topical overlaps misled the model, particularly in cases where evidence spanned multiple unrelated contexts or referenced similar entities in different time periods or locations.
Evidence quality and noise:
A.4 Automatic Filtering of Tulu3 prompts
Text:
Sentence to be focused on:
Facts:
Sentence to be focused on:
Facts:
Sentence to be focused on:
Sentence to be focused on:
Sentence to be focused on:
Facts:
No verifiable claim.
Text:
Sentence to be focused on:
Sentence to be focused on:
Facts:
No verifiable claim.
Input format:
### Prompt
<prompt here>
Output format:
### Thoughts
<your thoughts here>
### Judgement
Yes/No
@@ A.4 Automatic Filtering of Tulu3 prompts
Correct $\uparrow$ Incorrect $\downarrow$ Missing $\downarrow$
VeriScore & Tulu3 Personas, claim-level GPT-4o few-shot ICL 0.10 0.04 3.54 0.82 95.64 0.28
VeriFastScore-Stage-1 0.06 0.29 24.64 4.53 70.84 0.85
VeriFastScore 0.05 0.28 23.69 4.35 71.96 0.87
Tulu3 Personas, sentence-level GPT-4o few-shot ICL 0.15 0.03 2.61 0.57 96.82 0.33
VeriFastScore-Stage-1 0.06 0.19 15.33 3.43 81.24 0.75
VeriFastScore 0.05 0.23 18.62 4.18 77.26 0.80
Correct $\uparrow$ Incorrect $\downarrow$ Missing $\downarrow$
VeriScore data & Tulu3 Personas, claim-level GPT-4o few-shot ICL 0.30 0.31 18.1 7.2 73.2 0.28
VeriFastScore-Stage-1 0.88 0.87 63.0 21.3 15.1 0.85
VeriFastScore 0.87 0.90 71.6 15.5 12.3 0.87
Tulu3 Personas, sentence-level GPT-4o few-shot ICL 0.38 0.34 19.3 8.2 71.2 0.33
VeriFastScore-Stage-1 0.72 0.69 50.8 20.8 27.7 0.75
VeriFastScore 0.83 0.86 66 17.6 15 0.80
Table 10: Automatic performance metrics across test splits. VeriFastScore-Stage-1 refers to the best model trained with claim-level evidence. Evidence Granularity denotes whether evidence was retrieved using full sentences or claims. Precision and Recall measure overlap between claims identified by VeriFastScore and VeriScore. Under Claim Accuracy, Correct, Incorrect (Label), and Missing (Claim) capture agreement, label mismatches, and omissions relative to VeriScore. A small fraction of instances were marked Erroneous (not shown) due to ambiguities in automatic evaluation (see A.5). Correlation reports Pearson $r$ between factuality scores (see Eq. 1) assigned by VeriFastScore and VeriScore. For all Pearson’s correlation scores reported in this paper, p<0.001, unless stated otherwise.
A.5 Automatic Evaluation
To evaluate the factuality labels produced by VeriFastScore, we perform automatic evaluation using GPT-4o-mini as an entailment judge. Given an extracted claim and a set of reference claims (from VeriScore), the model is prompted to determine whether the claim can be inferred from the reference set. If supported, the model also identifies a minimal subset of reference claims that together justify the extracted claim. This setup ensures that factual consistency is assessed in a targeted, evidence-grounded manner while minimizing overreliance on superficial textual overlap.

### v3

# ΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥ VeriFastScore: Speeding up long-form factuality evaluation

@@ Abstract
ΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥΥ
VeriFastScore: Speeding up long-form factuality evaluation
Rishanth Rajendhran1   Amir Zadeh2   Matthew Sarte2   Chuan Li2   Mohit Iyyer1 1University of Maryland, College Park  2Lambda Labs 1{rishanth, miyyer}@umd.edu2{amirali.zadeh, matt.sarte, c}@lambdal.com
@@ 1 Introduction
[…]
@@ 2.3 VeriFastScore’s Integrated Approach
[…]
@@ 2.4 Generating Synthetic Training Data for VeriFastScore
We selected prompts from the Tulu3 Personas dataset Lambert et al. (2025), which contains diverse and high-quality instructions across domains intended to test model’s instruction-following ability and adaptation to user intent. This dataset was chosen because it aligns well with our goal of generating model responses that contain factual, verifiable content. We used GPT-4o-mini to identify a subset of prompts likely to elicit factual claims [^6] More details in A.3. A manual inspection of 200 filtered prompts showed that over 90% did indeed lead to factually dense responses.
@@
claim‐level GPT-4o few-shot ICL 0.30 0.31 18.1 7.2 73.2 0.28
@@
Table 3: Performance metrics across test splits: VeriFastScore produces claims and verification labels that are largely consistent with VeriScore, achieving a strong corelation of 0.80 even with sentence-level evidence, while significantly outperforming the GPT-4o few-shot prompting baseline. Evidence Granularity denotes whether evidence was retrieved using full sentences or claims: claim-level includes VeriScore data and Tulu3 Personas data while sentence-level includes Tulu3 Personas data only. Precision and Recall measure overlap between claims identified by VeriFastScore and VeriScore. Under Claim Accuracy, Correct, Incorrect (Label), and Missing (Claim) capture agreement, label mismatches, and omissions relative to VeriScore. A small fraction of instances were marked Erroneous (not shown) due to ambiguities in automatic evaluation (see A.2). Correlation reports Pearson $r$ between factuality scores (see Eq. 1) assigned by VeriFastScore and VeriScore. For all Pearson’s correlation scores reported in this paper, p<0.001, unless stated otherwise.
@@ 2.5 Evaluation Metrics
During early experiments, we found that when using exact match metrics, comparing model-extracted claims and verification labels directly to reference outputs, even minor surface-level differences (e.g., paraphrasing, slight reordering, or omission of non-essential modifiers) led to false negatives, thereby underestimating the model’s actual performance (see Appendix Table 8). For example, VeriFastScore achieved an exact match claim accuracy of 23.7% on the test set, versus 71.6% with paraphrase-aware evaluation. [^8] We re-report Table 8 with exact match metrics in Appendix.
@@ 3 Results
Response … Navigating a rainforest can pose several challenges … unpredictable weather which can lead to sudden rainstorms or flooding … VeriFastScore: (Supported) Unpredictable weather in a rainforest can lead to sudden rainstorms. VeriScore: (Unsupported) Navigating a rainforest can lead to sudden rainstorms.   In this example, VeriFastScore exhibits precise semantic alignment between extracted claim and model response while VeriScore does not.
@@ 3.2 Speedup
We measured the execution time for both the VeriFastScore and VeriScore pipelines. By retrieving evidence at a coarser granularity, VeriFastScore achieves a 39% reduction in API costs and cuts retrieval time by almost half. Additionally, VeriScore takes significantly longer due to its sequential claim extraction and verification processes. VeriFastScore is approximately 10 times faster in extracting and verifying claims without sacrificing performance (see Table 8). Overall, VeriFastScore achieves a 6.64x speedup compared to VeriScore. Longer model responses further increase speedup, as VeriScore’s sentence-by-sentence processing is particularly slow and expensive, in-part due to the larger number of SERPER queries VeriScore makes compared to VeriFastScore (see Table 6).
VeriScore 20.5 23 \$0.01725
VeriFastScore 58.8 14 \$0.0105
@@ 3.3 Ranking factuality of models with VeriFastScore
[…]
4 Runtime Analysis
To more precisely quantify the efficiency difference, we implemented a parallelized version of VeriScore and compared it directly with VeriFastScore. We distributed sentence-level claim decomposition and claim-level verification across 8 GPUs on a H100 node.
Parallelized VeriScore Implementation:
For a 100–instance subset of the test set, this setup reduced the claim-decomposition time to an average of 59.7s (compared to 253s for the standard single-GPU implementation), with a total modeling time (decomposition + verification) of approximately 235s.
Comparison with Unparallelized VeriFastScore:
By contrast, VeriFastScore processes the same 100 instances in only 29s on a single GPU on the same H100 node, while performing both claim extraction and verification in one pass.
Limits of Parallelization:
Although parallelization yields moderate speedups, VeriScore remains fundamentally constrained by its sequential pipeline: (1) claims must first be extracted sentence-by-sentence, and (2) each extracted claim must then be verified individually. These dependencies create repeated tokenization and input-formatting costs, along with inter-process communication overhead, all of which limit scalability and make efficient batching difficult. In contrast, VeriFastScore integrates decomposition and verification into a single forward pass over the entire response and consolidated evidence, eliminating intermediate stages and achieving consistent, end-to-end efficiency gains.
@@ 5 Qualitative Analysis
Claim grounding and hallucination:
All 401 extracted claims were grounded in the original model response, with no instances of hallucinated or fabricated claims. This is a crucial property for a factuality metric, particularly in downstream use cases like hallucination detection and RLHF training where spurious claim hallucination could mislead supervision.
Extraction coverage and omissions:
VeriFastScore was generally effective at identifying claims, especially in responses rich with factual content. However, it showed reduced reliability in responses that interweave factual and non-factual content, or contain few factual claims overall. Claim omission was observed in about 6 out of the 21 evaluated responses, though 4 of these cases involved fewer than four missing claims. Notably, in 3 of the omission cases, VeriFastScore extracted claims only from the final few sentences of the response, missing earlier factual content—suggesting a positional bias or early truncation.
Unverifiable and ill-formed claims:
About 42 out of the 401 extracted claims were judged to be unverifiable due to missing key details (e.g., named entities, dates) in the original model response. In rare cases, VeriFastScore also extracted claims from inherently unverifiable sentences, such as advice or vague generalizations. Ill-formed claims, those missing critical contextual elements, were rare (around 4 claims out of the total 401 claims), but they did alter the meaning of the original sentence in notable ways.
Verification accuracy:
VeriFastScore was largely accurate in labeling claims against the provided evidence, with an estimated verification error rate of 91.5%. Among the 34 incorrect verification labels, 25 of them were false positives: instances where VeriFastScore incorrectly marked a claim as supported. These errors often arose when partial matches or topical overlaps misled the model, particularly in cases where evidence spanned multiple unrelated contexts or referenced similar entities in different time periods or locations.
Evidence quality and noise:
@@ 7 Conclusion
[…]
@@ 8 Limitations
Our retrieval strategy issues Google search queries for each sentence in the model response, aggregating the results into a shared evidence context. While this aligns with the single-pass structure of VeriFastScore, it can introduce irrelevant or insufficient evidence, especially for vague or underspecified sentences. More adaptive or agentic retrieval, allowing the model to detect failures and issue improved follow-up queries, may yield higher evidence quality.
@@
VeriFastScore produces fine-grained factuality labels for claims but does not provide natural language justifications. Although rationale generation adds computational cost, it could improve interpretability and trustworthiness, especially in sensitive domains. Training with rationales, even if only used during finetuning (e.g., as in Qwen3’s “reasoning mode” Yang et al. (2025)), may offer benefits without incurring runtime overhead at inference time. Initial experiments on training a model to reason first using GRPO showed promise.
@@
We use GPT-4o-mini to assess label correctness and semantic equivalence, which enables scalable evaluation but introduces potential noise, particularly for nuanced paraphrase detection. This may affect metric estimates such as precision and recall, and obscure more subtle performance differences across models.
@@ A.1 Training VeriFastScore
As seen in Table 8, there was a $\sim$20% drop in claim accuracy when using sentence-level evidence, which is our true test setting, instead of claim-level evidence. This is possibly due to differences in quality of retrieved search results and length of consolidated evidence depending on the granularity at which evidence was collected. To mitigate this, the model was further trained on a mixture of claim- and sentence-level evidence in the second stage of training.
@@
A.2 Automatic Evaluation
To evaluate the factuality labels produced by VeriFastScore, we perform automatic evaluation using GPT-4o-mini as an entailment judge. Given an extracted claim and a set of reference claims, the model is prompted to determine whether the claim can be inferred from the reference set. If supported, the model also identifies a minimal subset of reference claims that together entail the extracted claim. More concretely, GPT-4o-mini was prompted to judge factual consistency of each predicted claim against gold claims, and vice versa. Factual precision is computed from predicted claims given the gold claims as the reference set, while factual recall is computed from gold claims given the predicted claims as the reference set. This two-way setup accounts for granularity mismatches (e.g., one composite claim vs. multiple atomic claims), which precludes one-to-one alignment while also ensuring that factual consistency is assessed in a targeted, evidence-grounded manner while minimizing overreliance on superficial textual overlap.
A.3 VeriFastScorer model
@@
Correct $\uparrow$ Incorrect $\downarrow$ Missing $\downarrow$
VeriScore & Tulu3 Personas, claim‐level GPT-4o few-shot ICL 0.10 0.04 3.54 0.82 95.64 0.28
VeriFastScore-Stage-1 0.06 0.29 24.64 4.53 70.84 0.85
VeriFastScore 0.05 0.28 23.69 4.35 71.96 0.87
Tulu3 Personas, sentence‐level GPT-4o few-shot ICL 0.15 0.03 2.61 0.57 96.82 0.33
VeriFastScore-Stage-1 0.06 0.19 15.33 3.43 81.24 0.75
VeriFastScore 0.05 0.23 18.62 4.18 77.26 0.80
Correct $\uparrow$ Incorrect $\downarrow$ Missing $\downarrow$
VeriScore data & Tulu3 Personas, claim‐level GPT-4o few-shot ICL 0.30 0.31 18.1 7.2 73.2 0.28
VeriFastScore-Stage-1 0.88 0.87 63.0 21.3 15.1 0.85
VeriFastScore 0.87 0.90 71.6 15.5 12.3 0.87
Tulu3 Personas, sentence‐level GPT-4o few-shot ICL 0.38 0.34 19.3 8.2 71.2 0.33
VeriFastScore-Stage-1 0.72 0.69 50.8 20.8 27.7 0.75
VeriFastScore 0.83 0.86 66 17.6 15 0.80
Table 9: Automatic performance metrics across test splits. VeriFastScore-Stage-1 refers to the best model trained with claim-level evidence. Evidence Granularity denotes whether evidence was retrieved using full sentences or claims. Precision and Recall measure overlap between claims identified by VeriFastScore and VeriScore. Under Claim Accuracy, Correct, Incorrect (Label), and Missing (Claim) capture agreement, label mismatches, and omissions relative to VeriScore. A small fraction of instances were marked Erroneous (not shown) due to ambiguities in automatic evaluation (see A.2). Correlation reports Pearson $r$ between factuality scores (see Eq. 1) assigned by VeriFastScore and VeriScore. For all Pearson’s correlation scores reported in this paper, p<0.001, unless stated otherwise.
@@
Text:
Sentence to be focused on:
Facts:
Sentence to be focused on:
Facts:
Sentence to be focused on:
Sentence to be focused on:
Sentence to be focused on:
Facts:
No verifiable claim.
Text:
Sentence to be focused on:
Sentence to be focused on:
Facts:
No verifiable claim.
Input format:
### Prompt
<prompt here>
Output format:
### Thoughts
<your thoughts here>
### Judgement
Yes/No